OpenStudy (anonymous):

What is the value of k? 3x + 7 = x2 + k = 7x + 15

OpenStudy (anonymous):

too many equal signs?

OpenStudy (anonymous):

That's the problem my teacher gave me.. ?

OpenStudy (anonymous):

okay, we can take the first and last equations 3x + 7 = 7x + 15 solve for x x = -2 and use this to solve for k

OpenStudy (anonymous):

Totally sure.

OpenStudy (anonymous):

Note 7x + 15 = 7(-2) + 15 = 1 Therefore 1 = x^2 + k or 1 = 4 + k so k = -3 QED

myininaya (myininaya):

3(-2)+7=(-2)^2+(-3)=7(-2)+15 1=1=1 fine maybe it is true lol

OpenStudy (anonymous):

Neat Question!

OpenStudy (anonymous):

Lol, thanks guys. ^_^